a positive charge of 3.6x10^-5 C and a negative charge of 2.4x10^-5 C are 0.034 meters apart calculate the magnitude of the electrostatic force between the two particles
1
Expert's answer
2020-04-02T10:38:57-0400
"F=\frac{kq_1q_2}{r^2}"

"F=\frac{(9\cdot 10^9)(3.6\cdot 10^{-5})(2.4\cdot 10^{-5})}{0.034^2}=6700\ N=6.7\ kN"